OpenStudy (anonymous):

what is the discontinuity and zero of the function f(x)=2x^2+5-12/x+4

OpenStudy (anonymous):

does your question look like this ? \[\frac{ 2x ^{2} + 5x -12 }{ x-4}\]

OpenStudy (anonymous):

ooops, x+4 in the denom

OpenStudy (anonymous):

@Stew.a.r.t. yes

OpenStudy (haseeb96):

http://www.math.brown.edu/UTRA/discontinuities.html read it it can help u

OpenStudy (anonymous):

im guessing the limit would be discontinuous for x tending towards -4, as that would give you a zero in the denominator any other value should work just fine

OpenStudy (anonymous):

should it be -4 zero at 3/2 or -3/2? @Stew.a.r.t.

OpenStudy (ranga):

At x = -4, the denominator goes to zero. Therefore, there is a discontinuity at x = -4. The function is not defined at x = -4. For finding the zeros, factor the numerator. See if any factor cancels out with the denominator. Then set whatever is remaining to zero and solve for x. (remember the domain of this function excludes x = -4 and therefore, the function cannot have a zero at x = -4).

OpenStudy (anonymous):

should it be -4 zero at 3/2 or -3/2? @ranga

OpenStudy (ranga):

Discontinuity at x = -4 Zero at x = 3/2

OpenStudy (ranga):

The function does NOT go to zero at x = -3/2. It does for x = 3/2.

OpenStudy (anonymous):

ok thank you

OpenStudy (ranga):

yw.